OpenStudy (anonymous):

5L/min go into a container and 2.4L/min come out with an initial volume of 10L how full is the container in 5min?

OpenStudy (anonymous):

10 + 5*5 - 2.4 * 5 10 + 25 - 12 23

OpenStudy (anonymous):

excelent!

OpenStudy (amistre64):

How much is being kept in each min? Total = in - out Total = 5 - 2.4 = 2.6 2.6 l stays in each minute after 5 min: 5(2.6) = 13 L are retained after 5 min. If the initial volume is 10: 10 + 13 = 23 ....

OpenStudy (anonymous):

Bcome a fan!